one number is 6 more than the other. their sum is o. what is the value of the smaller number?: one number is 6 more than the other. their sum is o. what is the value of the smaller number?
Answer by jim_thompson5910(8743)

"one number is 6 more than the other" translates to y=x+6
"Their sum is 0" translates to x+y=0



x+x+6=0 Plug in y=x+6 into the second equation


2x+6=0 Combine like terms on the left side.


2x=0-6 Subtract 6 from both sides.


2x=-6 Combine like terms on the right side.


x=(-6)/(2) Divide both sides by 2 to isolate x.


x=-3 Reduce.


y=x+6 Go back to the first equation


y=-3+6 Plug in x=-3


y=3 Simplify


So the answer is x=-3 and y=3